# Style Content Decomposition-based Data Augmentation for Domain Generalizable Medical Image Segmentation

### v2

@@ Introduction
We devise image content augmentation to mitigate content shifts (local disparities in anatomical structure). This is achieved by mixing a set of $t$ randomly selected content maps (i.e., rank-one matrices) from an image $x_{i}$ with another set of $t$ randomly selected content maps from an auxiliary image $x_{j}$ [Fig. 2]. The perturbation to each content map is formulated as:
$\bar{\boldsymbol{u}}_{i}^{k}=\beta\times\boldsymbol{u}_{i}^{k}+(1-\beta)\times\boldsymbol{u}_{j}^{k}\quad or\quad\bar{\boldsymbol{v}}_{i}^{k}=\beta\times\boldsymbol{v}_{i}^{k}+(1-\beta)\times\boldsymbol{v}_{j}^{k}$ (3)
@@
where the weight $\beta\sim U(0,1)$ and $\boldsymbol{u}_{i/j}^{k}$/$\boldsymbol{v}_{i/j}^{k}$ represents the left/right singular vector of image $x_{i/j}$. Using the perturbed content maps and the original style codes, we can obtain a content-augmented image $\bar{x}_{i}$ via Eq. (1).
@@
In a nutshell, by leveraging Eq. (1)(2)(3), StyCona conducts style and content augmentation and produces an augmented image $\bar{\tilde{x}}_{i}=\sum^{r_{i}}_{k=1}\tilde{\boldsymbol{\sigma}}_{i}^{k}\bar{\boldsymbol{u}}_{i}^{k}{\bar{\boldsymbol{v}}_{i}^{k}}{}^{\mathsf{T}}$.
Loss Supervision
@@
Afterwards, we formulate the segmentation loss $\mathcal{L}$ for training a domain generalizable model $f(\cdot;\theta)$ based on StyCona images:
@@
The MS-CMR Dataset (Single Source and Single Target Domain) dataset (Zhuang 2018) contains 45 subjects, each with bSSFP and LGE MRI sequences, along with ground truth annotations for the right ventricle (RV), left ventricle (LV), and myocardium (MYO). All images are normalized to the range [0,1][0,1][ 0 , 1 ] and resampled to a uniform resolution of $\times 1.0$ 1.0 mm. This experimental setting evaluates a segmentation model’s ability to generalize across MRI sequences for cardiac structure segmentation, specifically from LGE to bSSFP and vice versa. For each direction, one sequence (e.g., bSSFP) is treated as the source domain and is divided into training and validation sets with a 8:28:28 : 2 ratio, while another sequence (e.g., LGE) serves as the target domain and is used solely for testing.
@@
The Fundus Image Benchmark (Single Source and Multiple Target Domains(Chen et al. 2023) is established based on the ORIGA (Zhang et al. 2010), Drishti-GS (Sivaswamy et al. 2014), REFUGE (Orlando et al. 2020), and RIGA (Almazroa et al. 2018) (BinRushed and Magrabia) datasets. This setting evaluates the model’s generalization capability for the joint OC and OD segmentation across multiple target domains. The ORIGA dataset is used as the source domain with the training and validation split following TriD (Chen et al. 2023); the other datasets (BinRushed, Drishti-GS, Magrabia, and REFUGE) serve as target domains for cross-domain evaluation.
@@
We conducted the experiments using PyTorch (Paszke et al. 2019) on an NVIDIA A40 GPU with 48G GPU memory. All compared methods were optimized using an AdamW optimizer (Kingma and Ba 2015) with a fixed learning rate of $1e-4$ during the 100 training epochs for both the cadiac MRI and fundus image segmentation tasks. U-Net (Ronneberger, Fischer, and Brox 2015) was employed as the baseline segmentation model. The combination of cross-entropy and Dice loss (Milletari, Navab, and Ahmadi 2016) was used as the segmentation criterion. All images were resized to $\times 256$ 256 for both training and testing. During evaluation, the predicted labels were rescaled to their original resolutions. We set the number of perturbed content maps $t=16$ (please refer to the ablation study for further analysis).
Comparison with State of the Arts
We extensively compared our StyCona with state-of-the-art DG methods: 1) Fourier transformation-based: amplitude swap (AmpSwap) (Yang and Soatto 2020), amplitude mixup (AmpMix) (Xu et al. 2021a), and FMAug (Li et al. 2023)); 2) Feature statistics editing-based: MixStyle (Zhou et al. 2021), TriD (Chen et al. 2023), and ConStyX (Chen et al. 2025); 3) Random convolution-based: RandConv (Xu et al. 2021b), CIDA (Ouyang et al. 2022), and PRandConv (Choi et al. 2023). We re-implemented all the compared methods in a unified experimental setup for fair comparison. All the methods are built upon the same baseline segmentation model (U-Net). In general, StyCona sets the state-of-the-art in all three settings, both quantitatively [Table 1 and Table 2] and qualitatively [Fig. 3], showcasing its superiority in domain generalizable medical image segmentation.
@@
U-Net (Ronneberger, Fischer, and Brox 2015) 65.97 6.55 76.91 3.41
AmpSwap (Yang and Soatto 2020) 70.93 6.48 78.74 3.26
AmpMix (Xu et al. 2021a) 63.89 6.44 80.07 2.46
FMAug (Li et al. 2023) 72.46 13.67 80.54 2.57
MixStyle (Zhou et al. 2021) 69.26 10.91 80.81 2.69
TriD (Chen et al. 2023) 65.12 10.25 79.14 4.60
ConStyX (Chen et al. 2025) 68.63 7.89 81.30 2.85
RandConv (Xu et al. 2021b) 71.90 5.65 75.17 4.27
CIDA (Ouyang et al. 2022) 69.72 8.63 72.85 4.61
PRandConv (Choi et al. 2023) 56.61 8.75 75.57 3.38
@@
Table 2: Comparison with state-of-the-art methods on the fundus segmentation tasks with 10% labeled data (Source: ORIGA; Target: REFUGE, Drishti-GS, BinRushed, and Magrabia). The best and second-best results are highlighted in bold and underline, respectively.
@@
U-Net (Ronneberger, Fischer, and Brox 2015) 54.82 7.07 77.99 1.16 61.19 8.10 79.49 1.89 68.37 4.56
AmpSwap (Yang and Soatto 2020) 63.56 4.36 78.16 1.13 66.13 4.45 81.84 1.45 72.41 2.85
AmpMix (Xu et al. 2021a) 65.08 4.32 77.06 1.09 68.14 4.29 79.50 1.67 72.44 2.84
FMAug (Li et al. 2023) 63.79 5.03 78.34 1.04 64.33 7.28 82.17 1.29 72.16 3.66
MixStyle (Zhou et al. 2021) 63.95 4.72 80.36 1.06 63.97 6.04 81.93 1.45 72.55 3.32
TriD (Chen et al. 2023) 66.27 4.34 77.08 1.24 65.42 4.53 79.36 1.57 72.03 2.92
ConStyX (Chen et al. 2025) 62.91 5.99 80.69 1.03 64.47 7.16 80.15 1.85 72.05 4.01
RandConv (Xu et al. 2021b) 63.75 4.70 79.35 1.13 65.61 5.07 81.52 1.46 72.56 3.09
CIDA (Ouyang et al. 2022) 53.16 11.34 78.55 5.18 51.27 14.61 68.09 6.62 62.76 8.44
PRandConv (Choi et al. 2023) 44.60 9.84 75.23 1.48 53.34 8.68 76.61 2.57 62.44 5.64
@@
Results on Cardiac MRI
As reported in Table 1, StyCona achieves consistent performance improvements over the compared methods in the cardiac MRI segmentation task. In general, almost all the compared methods surpass the baseline model, indicating their effectiveness in addressing domain shifts. However, it is worth noting that some augmentation methods, e.g., AmpMix (Xu et al. 2021a) and TriD (Chen et al. 2023), yield segmentation results inferior to those of the baseline model in the bSSFP $\rightarrow$ LGE scenario. This may stem from the fact that the blended amplitude spectrum in AmpMix and the mixed feature statistics in TriD generate style-augmented images that are insufficient for mitigating content shifts, resulting in the models overfitting spurious correlations between the augmented images and their corresponding segmentation labels. In contrast, StyCona demonstrates a clear advantage in handling both style and content shifts by style content augmentation. These results validate the core idea of StyCona and highlight its effectiveness in domain-generalizable medical image segmentation.
Results on OC/OD Fundus Image
In this experiment, we further evaluate StyCona in a single-source, multi-target domain setting. As shown in Table 2, all methods generally achieve comparable average performance across the four target domains. However, the two random convolution-based approaches (i.e., CIDA (Ouyang et al. 2022) and PRandConv (Choi et al. 2023)) yield unsatisfactory results. This may be attributed to that the random convolution operations produce augmented images with distorted content that is inconsistent with the corresponding segmentation labels, thereby misleading the segmentation models during training. In contrast, StyCona demonstrates more robust performance across the four target domains, achieving the highest average DSC and competitive ASD. These results further suggest the effectiveness of StyCona in domain-generalizable medical image segmentation with multiple target domains.
Qualitative results.
As shown in Fig. 3, StyCona produces segmentation results with more precise object delineation. The improvement can be attributed to StyCona’s style and content augmentations, which enable the augmented samples to cover a wide range of unseen domains, thereby allowing the trained model to generate accurate segmentation maps for unseen data. In comparison, methods such as AM (Xu et al. 2021a), MixStyle (Zhou et al. 2021), and PRandConv (Choi et al. 2023) yield relatively unsatisfactory results. AM and MixStyle rely solely on style augmentation by blending amplitude spectra and adjusting the mean and standard deviation of samples, respectively. Meanwhile, the effectiveness of PRandConv is limited by its progressively applied random convolutions, which generate content-distorted images and fail to adequately alleviate variations in anatomical structures. These qualitative results, consistent with the quantitative performance, further suggest the effectiveness of our approach.
@@
Table 3: Ablation study of the proposed method on the cross-sequence setting. The best results are highlighted in bold.
@@
Ablation Study
Effectiveness of each component
In Table 3, we present ablation experiments to analyze the contributions of each component of StyCona. These experiments were conducted under the cardiac MRI segmentation task. The results reveal a consistent trend: segmentation performance improves as the proposed style and content augmentation components are progressively integrated into our method. Specifically, the baseline model (i.e., U-Net) performs relatively unsatisfactorily on the target domain, as it tends to learn domain-specific decision rules based on source domain data. Then, the segmentation performance increases by a large gap when the style augmentation operation is introduced, effectively mitigating style shifts, which are an essential part of domain shifts in these scenarios. In contrast, the improvement is more significant, with an increase of over 7% in DSC, when only the content augmentation component is introduced to address the content shifts, another key aspect of domain shifts. Finally, by incorporating both the style and content augmentation modules to address global variations in image appearance and local differences in anatomical structures, our final method achieves the best performance with DSC of 73.39% and 81.59% on the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios, respectively. This result demonstrates the effectiveness of each component and validates our assumption that style and content augmentations can alleviate the two types of domain shifts.
Impact of No. Perturbed Content Maps
According to Eq. (1)(2)(3), perturbing the style codes alters only an image’s style without affecting its target shape corresponding to segmentation labels. On the other hand, perturbing $t$ randomly selected content maps results in changes in local anatomical structures (e.g., background tissues or target object-related boundaries), which is the goal of our method to enhance the robustness of segmentation models to anatomical structure variations. Qualitatively, Fig. 4(a) shows the StyCona-augmented samples with varying numbers of perturbed content maps ($t=8,16,32$). Compared with the original images, the augmented images with $t=8,16$ exhibit variations in overall image appearance and regional anatomical structures while remaining consistent with the ground truth segmentation labels. In contrast, at $t=32$, excessive perturbation is introduced into some target boundaries, which may alter their ground truth labels. Quantitatively, among all choices, $t=16$ yields the highest DSC in both the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios. Based on these results, we set $t=16$ in StyCona to balance perturbation strength with the invariance of segmentation labels, ensuring label consistency after content augmentation.
Conclusion
We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation. Our findings reveal that an image’s singular values function as style codes, governing its global image properties, while its rank-one matrices serve as content maps, determining its anatomical structures. Based on this decomposition, we categorize domain shifts into style and content shifts w.r.t. deviations in style codes and content maps, respectively. To address these shifts, StyCona perturbs style codes for style augmentation and blends content maps for content augmentation. Extensive experiments on cardiac MRI segmentation with a single target domain and OC/OD fundus image segmentation with multiple target domains demonstrate that StyCona is a promising data augmentation technique for domain-generalizable medical image segmentation.

### v3

@@ 2.1.3 Content Augmentation
We devise content augmentation to mitigate content shifts (local disparities in anatomical structure). This is achieved by mixing a set of $t$ randomly selected content maps from an image $x_{i}$ with another set of $t$ randomly selected content maps from an auxiliary image $x_{j}$ [Fig. 2]. The perturbation to each content map is formulated as:
$\bar{\boldsymbol{u}}_{i}^{r}=\beta\times\boldsymbol{u}_{i}^{r}+(1-\beta)\times\boldsymbol{u}_{j}^{r}\quad or\quad\bar{\boldsymbol{v}}_{i}^{r}=\beta\times\boldsymbol{v}_{i}^{r}+(1-\beta)\times\boldsymbol{v}_{j}^{r}$ (3)
@@
where the weight $\beta\sim U(0,1)$ and $\boldsymbol{u}_{i/j}^{r}$$\boldsymbol{v}_{i/j}^{r}$ represents the left/right singular vector of image $x_{i/j}$.
@@
In a nutshell, StyCona performs style and content augmentation by Eq. (1)(2)(3), and generates an augmented image as $\bar{\tilde{x}}_{i}=\sum^{R}_{r=1}\tilde{\boldsymbol{\sigma}}_{i}^{r}\bar{\boldsymbol{u}}_{i}^{r}{\bar{\boldsymbol{v}}_{i}^{r}}{}^{\mathsf{T}}$.
2.1.4 Loss Supervision
@@
Afterwards, we formulate the segmentation loss $\mathcal{L}$ for training a domain generalizable model $f(\cdot;\theta)$ based on StyCona augmented images:
@@ 3 Experiments and Results
The MS-CMR Dataset (Single Source and Single Target Domain) dataset (zhuang2018multivariate) contains 45 subjects, each with bSSFP and LGE MRI sequences, along with ground truth annotations for the right ventricle (RV), left ventricle (LV), and myocardium (MYO). All images are normalized to the range $[0,1]$ and resampled to a uniform resolution of $1.0\times 1.0$ mm. This experimental setting evaluates a segmentation model’s ability to generalize across MRI sequences for cardiac structure segmentation, specifically from LGE to bSSFP and vice versa. For each direction, one sequence (e.g., bSSFP) is treated as the source domain and is divided into training and validation sets with a $8:2$ ratio, while another sequence (e.g., LGE) serves as the target domain and is used solely for testing.
@@
The Fundus Image Benchmark (Single Source and Multiple Target Domains(chen2023treasure) is established based on the ORIGA (zhang2010origa), Drishti-GS (sivaswamy2014drishti), REFUGE (orlando2020refuge), and RIGA (almazroa2018retinal) (BinRushed and Magrabia) datasets. This setting evaluates the model’s generalization capability for the joint OC and OD segmentation across multiple target domains. The ORIGA dataset is used as the source domain with the training and validation split following TriD (chen2023treasure); the other datasets (BinRushed, Drishti-GS, Magrabia, and REFUGE) act as target domains for cross-domain evaluation.
@@
We conducted the experiments using PyTorch (paszke2019pytorch) on an NVIDIA A40 GPU with 48G GPU memory. All compared methods were optimized using an AdamW optimizer (kingma2014adam) with a fixed learning rate of 1e41e-4 during the 100 training epochs for both the cadiac MRI and fundus image segmentation tasks. U-Net (ronneberger2015u) was employed as the baseline segmentation model. The combination of cross-entropy and Dice loss (milletari2016v) was used as the segmentation criterion. All images were resized to $256\times 256$ for both training and testing, and the predicted labels were rescaled to their original resolutions for evaluation. We set the number of perturbed content maps $t=16$ (please refer to Section 3.2.2 for further analysis).
3.1 Comparison with State of the Arts
We compared StyCona with state-of-the-art DG methods: 1) Fourier transformation-based: amplitude swap (AmpSwap) (yang2020fda), amplitude mixup (AmpMix) (xu2021fourier), and FMAug (li2023frequency)); 2) Feature statistics editing-based: MixStyle (zhou2021mixstyle), TriD (chen2023treasure), and ConStyX (chen2025constyx); 3) Random convolution-based: RandConv (xu2021robust), CIDA (ouyang2022causality), and PRandConv (choi2023progressive). We re-implemented all the compared methods in a unified experimental setup for fair comparison. All the methods are built upon the same baseline segmentation model (U-Net). In general, StyCona sets the state-of-the-art in all three settings, both quantitatively [Table 1 and Table 2] and qualitatively [Fig. 3], showcasing its superiority in domain generalizable medical image segmentation.
3.1.1 Results on Cardiac MRI
As reported in Table 1, StyCona achieves consistent performance improvements over the compared methods in the cardiac MRI segmentation task. In general, almost all the compared methods surpass the baseline model, indicating their effectiveness in addressing domain shifts. However, it is worth noting that some methods, e.g., AmpMix (xu2021fourier) and TriD (chen2023treasure), yield segmentation results inferior to those of the baseline model in the bSSFP $\rightarrow$ LGE scenario. This may result from the fact that the blended amplitude spectrum in AmpMix and the mixed feature statistics in TriD generate style-augmented images that are insufficient for mitigating content shifts, resulting in the models overfitting spurious correlations between the augmented images and their corresponding segmentation labels. In contrast, StyCona demonstrates a clear advantage in handling both style and content shifts by style content augmentation. These results validate the key idea of StyCona and highlight its effectiveness in domain generalizable medical image segmentation.
3.1.2 Results on Fundus Image
We further evaluate StyCona in a single-source, multi-target domain setting. As shown in Table 2, all methods generally achieve comparable average performance across the four target domains. However, the two random convolution-based approaches (i.e., CIDA (ouyang2022causality) and PRandConv (choi2023progressive)) yield relatively unsatisfactory results. This may be attributed to that the random convolution operations produce augmented images with distorted content that is inconsistent with the corresponding segmentation labels, thus misleading the segmentation models during training. In contrast, StyCona shows more robust performance across the four target domains, obtaining the highest average DSC and competitive ASD. These results further suggest the effectiveness of StyCona in domain generalizable medical image segmentation with multiple target domains.
@@ 3.1.2 Results on Fundus Image
U-Net (ronneberger2015u) 65.97 6.55 76.91 3.41
AmpSwap (yang2020fda) 70.93 6.48 78.74 3.26
AmpMix (xu2021fourier) 63.89 6.44 80.07 2.46
FMAug (li2023frequency) 72.46 13.67 80.54 2.57
MixStyle (zhou2021mixstyle) 69.26 10.91 80.81 2.69
TriD (chen2023treasure) 65.12 10.25 79.14 4.60
ConStyX (chen2025constyx) 68.63 7.89 81.30 2.85
RandConv (xu2021robust) 71.90 5.65 75.17 4.27
CIDA (ouyang2022causality) 69.72 8.63 72.85 4.61
PRandConv (choi2023progressive) 56.61 8.75 75.57 3.38
@@
Table 2: Comparison with state-of-the-art methods on the fundus image segmentation tasks (Source: ORIGA; Target: REFUGE, Drishti-GS, BinRushed, and Magrabia). The best and second-best results are highlighted in bold and underline, respectively.
@@
U-Net (ronneberger2015u) 54.82 7.07 77.99 1.16 61.19 8.10 79.49 1.89 68.37 4.56
AmpSwap (yang2020fda) 63.56 4.36 78.16 1.13 66.13 4.45 81.84 1.45 72.41 2.85
AmpMix (xu2021fourier) 65.08 4.32 77.06 1.09 68.14 4.29 79.50 1.67 72.44 2.84
FMAug (li2023frequency) 63.79 5.03 78.34 1.04 64.33 7.28 82.17 1.29 72.16 3.66
MixStyle (zhou2021mixstyle) 63.95 4.72 80.36 1.06 63.97 6.04 81.93 1.45 72.55 3.32
TriD (chen2023treasure) 66.27 4.34 77.08 1.24 65.42 4.53 79.36 1.57 72.03 2.92
ConStyX (chen2025constyx) 62.91 5.99 80.69 1.03 64.47 7.16 80.15 1.85 72.05 4.01
RandConv (xu2021robust) 63.75 4.70 79.35 1.13 65.61 5.07 81.52 1.46 72.56 3.09
CIDA (ouyang2022causality) 53.16 11.34 78.55 5.18 51.27 14.61 68.09 6.62 62.76 8.44
PRandConv (choi2023progressive) 44.60 9.84 75.23 1.48 53.34 8.68 76.61 2.57 62.44 5.64
@@
3.1.3 Qualitative results.
As shown in Fig. 3, StyCona produces segmentation results with more precise object delineation. The improvement can be attributed to StyCona’s style and content augmentations, which enable the augmented samples to cover a wide range of unseen domains, thereby allowing the trained model to generate accurate segmentation maps for unseen data. In comparison, methods such as AmpMix (xu2021fourier), MixStyle (zhou2021mixstyle), and PRandConv (choi2023progressive) yield relatively unsatisfactory results. AmpMix and MixStyle rely solely on style augmentation by blending amplitude spectra and adjusting the mean and standard deviation of samples, respectively. Meanwhile, the effectiveness of PRandConv is limited by its progressively applied random convolutions, which generate content-distorted images and fail to adequately alleviate variations in anatomical structures. These qualitative results, consistent with the quantitative performance, further suggest the effectiveness of our approach.
@@
3.2 Ablation Study
3.2.1 Effectiveness of each component
In Table 3, we present ablation experiments to analyze the contributions of each component of StyCona. These experiments were conducted under the cardiac MRI segmentation task. The results reveal a consistent trend: segmentation performance improves as the proposed style and content augmentation components are gradually integrated into our method. Specifically, the baseline model (i.e., U-Net) performs relatively unsatisfactorily on the target domain, as it tends to learn domain-specific decision rules based on source domain data. Then, the segmentation performance increases by a large gap when the style augmentation operation is introduced for effectively mitigating style shifts, which are an essential part of domain shifts in these scenarios. In contrast, the improvement is more significant, with an increase of over 7% in DSC, when only the content augmentation component is introduced to address the content shifts, another key aspect of domain shifts. Finally, by incorporating both the style and content augmentation modules to address global variations in image appearance and local differences in anatomical structures, our final method achieves the best performance with DSC of 73.39% and 81.59% on the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios, respectively. This result demonstrates the effectiveness of each component and validates our assumptions of the style content decomposition and the corresponding style and content augmentation to alleviate the domain shifts.
3.2.2 Impact of No. Perturbed Content Maps
According to Eq. (1)(2)(3), perturbing the style codes alters only an image’s style without changing its semantic information corresponding to segmentation labels. On the other hand, perturbing $t$ randomly selected content maps results in changes in local anatomical structures (e.g., background tissues or target object-related boundaries), which is the goal of StyCona to enhance the robustness of segmentation models to local anatomical structure variations. Qualitatively, Fig. 4(a) shows the StyCona-augmented samples with varying numbers of perturbed content maps ($t=8,16,32$). Compared with the original images, the augmented images with $t=8,16$ exhibit variations in global image appearance and local anatomical structures while remaining consistent with the ground truth segmentation labels. In contrast, excessive perturbation is introduced into some target boundaries when $t=32$, which may alter their ground truth labels. Quantitatively, $t=16$ yields the highest DSC in both the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios. Based on these results, we set $t=16$ in StyCona to balance perturbation strength with the invariance of segmentation labels, ensuring label consistency after content augmentation.
@@ 3.2.1 Effectiveness of each component
Table 3: Ablation study of StyCona on the cross-sequence setting. The best results are highlighted in bold.
@@
4 Conclusion
We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation. Our style content decomposition strategy reveals that an image’s singular values function as style codes, governing its global image properties, while its rank-one matrices serve as content maps, determining its anatomical structures. Based on this decomposition, we categorize domain shifts into style and content shifts w.r.t. deviations in style codes and content maps, respectively. Then, StyCona perturbs style codes for style augmentation and blends content maps for content augmentation, addressing both types of domain shifts accordingly. Extensive experiments on cardiac MRI segmentation with a single target domain and fundus image segmentation with multiple target domains demonstrate that StyCona is an effective data augmentation technique for domain-generalizable medical image segmentation.